OpenStudy (anonymous):

index4 sqrt 10,000 * index4 sqrt 100,00

jimthompson5910 (jim_thompson5910):

\[\large \sqrt[4]{10000}\times\sqrt[4]{10000}\] \[\large \sqrt[4]{10000\times10000}\] \[\large \sqrt[4]{100000000}\] \[\large 100\] So \[\large \sqrt[4]{10000}\times\sqrt[4]{10000}=100\]